OpenStudy (anonymous):

Find the product of the squre root of 30 and 6 to the sq root of 10 could someone help me please

OpenStudy (anonymous):

\[\sqrt{30} * \sqrt{6} = \sqrt{2*3*5} * \sqrt{2*3}\] This can be written \[\sqrt{2*2*3*3*5}\] This can be written \[\sqrt{4}*\sqrt{9}*\sqrt{5}\] Which can be written 2*3\[\sqrt{5}\] Which is 6\[\sqrt{5}\]

OpenStudy (anonymous):

Now we multiply the previous answer by \[\sqrt{10}\] Again write out fully = 6*\[\sqrt{5}*\sqrt{2}*\sqrt{5}\] =6*5*\[\sqrt{2}\] = 30 * \[\sqrt{2}\]